OpenStudy (anonymous):

-What are the explicit equation and domain for a geometric sequence with a first term of 3 and a second term of −9? an = 3(−12)n − 1; all integers where n ≥ 1 an = 3(−12)n − 1; all integers where n ≥ 0 an = 3(−3)n − 1; all integers where n ≥ 1 an = 3(−3)n − 1; all integers where n ≥ 0

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

please help

OpenStudy (campbell_st):

well with a 1st term 3 the common ratio would appear to be r = -3 becuase 3 x -3 = -9 the standard from of the equation is \[a_{n} = a_{1} \times r^{n - 1}\] so the 1st term is the clue.... 3 1st means n >= 1 and you know r = -3 so substituting you get \[a_{n} = 3 \times(-3)^{n - 1}\] hope it helps

OpenStudy (anonymous):

so if you had to pick a answer which one would it be @campbell_st

OpenStudy (campbell_st):

well I thought my answer was rather obvious as to which is the correct answer.

OpenStudy (anonymous):

@campbell_st so it would most likely be C

OpenStudy (campbell_st):

that's most likely right

OpenStudy (anonymous):

ok @campbell_st thanks